arielle is 1 1/3 times as old as pete. 5 years ago, her age was 1 1/2 his. Find arielle's present age
Answer by jorel1380(3719)

Arielle equals 4/3 Pete
Arielle minus 5 equals 3/2(Pete-5)
4/3P-5=3/2Pete-15/2
10/2+15/2=(9/6-8/6)P
5/2=P/6
2P=30
Pete=15
Arielle=20.